OpenStudy (anonymous):

Solve this for me please? I dont understand the quadratic formula. 3x^2 = 33x + 24

OpenStudy (anonymous):

x=-2.667 x=-1

OpenStudy (anonymous):

the quantity of negative 11 plus or minus 3 square root 89 all over 2 the quantity of 11 plus or minus square root 153 all over 2 the quantity of 11 plus or minus square root 89 all over 2 the quantity of negative 11 plus or minus 9 square root 17 all over 2 these are the options

OpenStudy (anonymous):

goodpoint I dont understand these answer choices but i know the formula

OpenStudy (anonymous):

x=(-33 plus or minus square root of 33-4*-3*24)over(2*-3)

OpenStudy (anonymous):

- 3x² + 33x + 24 = 0 Roots: 11.68465843, -0.6846584384 Discriminant: 1377 Vertex: (5.5, 114.75) Factored: f(x) = -3(x - 11.68465843)(x + 0.6846584384) Sum of Roots (-b/a): 11 Product of Roots (c/a): -8

OpenStudy (anonymous):

https://www.desmos.com/calculator/165212a7ba

jhonyy9 (jhonyy9):

3(x2 -11x-8)=0 x2 -11x -8 =0 x_1,_2=(11+/- sqrt(121+32))/2 =(11+/- sqrt(153))/2 =(11+/-3sqrt17)/2